82. Driving marathon. Felix drove 800 miles in x hours on
Monday.
a) Write a rational expression for his average speed.
b) On Tuesday he drove for 6 hours at the same average
speed. Write a rational expression for his distance on
Tuesday.

Answer by stanbon(75887)

Felix drove 800 miles in x hours on
Monday.
a) Write a rational expression for his average speed.
rate = d/t
rate = 800/x mph
------------------------
b) On Tuesday he drove for 6 hours at the same average
speed. Write a rational expression for his distance on
Tuesday.
d = rt
d = (800/x)6
d = 4800/x miles
======================
Cheers,
Stan H.